1) When 0.519 mol of HCOOH is dissolved in 0.409 L of water, proton transfer occurs.

HCOOH + H2O ⇄HCOO- + H3O+

The equilibrium concentration of H3O+ ions is 0.0149 M. Evaluate Keq.
Answer #1

Initial concentration of HCOOH = mol of HCOOH / volume in L

= 0.519 mol / 0.409 L

= 1.27 M

HCOOH + H2O <-> HCOO- + H3O+

1.27 0 0 (initial)

1.27-x x x (at equilibrium)

Given at equilibrium,

[H3O+] = 0.0149 M

So,

x = 0.0149 M

Use:

Keq = [HCOO-][H3O+] / [HCOOH]

= x*x / (1.27-x)

= (0.0149* 0.0149) / (1.27-0.0149)

= 1.77*10^-4

Answer: 1.77*10^-4